OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2). y = -one thirdx - 6 y = -one thirdx - 14 y = 3x - 6 y = 3x - 14

OpenStudy (solomonzelman):

\[3x - y = -3\]\[ - y = -3-3x\]\[y=3x+3\]this is the line. parallel, but passes through (4,-2) the slope will be same, but the y-intercept will change. \[y=3x+b\]\[(4)=3(-2)b\]\[4=-6+b\]\[10=b\]\[y=3x+10\]There.

OpenStudy (solomonzelman):

ops the other way.

OpenStudy (solomonzelman):

\[(-2)=(4)3+b\]

OpenStudy (solomonzelman):

\[-2=12+b\]\[-14=b\]\[y=3x-14\]\[IT'S~~~~D\]